OpenStudy (anonymous):

derivative of [ln(x)]^2 ;x=e^2

OpenStudy (wikiemol):

e^2 is just a constant so the whole derivative would be 0.

hartnn (hartnn):

you need to use chain rule, know what it is ?

OpenStudy (anonymous):

U gotta use the property of natural logs as well as the chain rule. so u get: 2(ln(x))*1/x for e^2 its the rule of e, which says that e^x is equal to e^x, so e^2 = e^2

OpenStudy (anonymous):

The chain rule would be the derivative 2[ln(x)]^1 times (1/x),right?

hartnn (hartnn):

that is correct, now just plug in x=e^2

OpenStudy (anonymous):

oh u had to plug in? in that case, ignore the second part of my answer.

OpenStudy (anonymous):

2ln(e^2)*(1/e^2) would simplify to 2e/e^2 to 2/e?

hartnn (hartnn):

ln e=1 so just 2/e^2

hartnn (hartnn):

sorry, 4/e^2

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

so dont the ln and e cancel, leaving behind an e?

OpenStudy (anonymous):

sorry, the ln and e^2 cancel...right?

OpenStudy (anonymous):

I just don't understand this canceling part~

OpenStudy (anonymous):

it should be 4/e

OpenStudy (anonymous):

No, hartnn's answer is right, and I have the answers, I just dont know how to get to it~

OpenStudy (anonymous):

2 * 1/e^2 * 2e

OpenStudy (anonymous):

how did ln(e^2) become 2e?

OpenStudy (anonymous):

i am using the chain rule

OpenStudy (anonymous):

I used it and came out with 2*ln(e^2)*(1/e^2) but I dont know how to go from there

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

u didnt take the derivate of e^2 yet

OpenStudy (anonymous):

U just substitute it right in, right...?

OpenStudy (anonymous):

substitute?

OpenStudy (anonymous):

yeah, x=e^2, where i plugged it in when I couldn't go any farther with the x.

OpenStudy (anonymous):

so what's ln(e^2) simplified?

OpenStudy (anonymous):

1/e^2 * 2e

hartnn (hartnn):

use the property \(\large \ln a^b=b\ln a\) so, \(\ln e^2 = 2\ln e\) and ln e = 1

hartnn (hartnn):

\(\ln e^2=2\)

OpenStudy (anonymous):

no, not talking about the deriavtive of ln(e^2), Nageer ,simplified form~

hartnn (hartnn):

(2 ln e^2)/e^2= (2*2)/e^2 = 4/e^2

OpenStudy (anonymous):

OHHH I get it now, hartnn... but why wouldn't it be e? cos ln(e^2) is the same as ln(e)(e), right? OH, you can't take the log of one and not the other... oops :D got it

hartnn (hartnn):

sorry for the delayed reply...

OpenStudy (anonymous):

itz ok, u were offline for a little~ :)

OpenStudy (anonymous):

well, thanx!!! I got it~! well then, hasta luego!

hartnn (hartnn):

welcome ^_^